solve for 4/9 = x/x3, and what is the formula to solve for this?
Answer by user_dude2008(1825)

4/9 = x/x^3

4/9 = 1/x^2

4x^2=9

x^2=9/4

x=3/2 or x=-3/2 <---------- Answers